Calculate the GDP using EXPENDITURE APPROACH with the following hypothetical data:





Personal Consumption 3657





Depreciation 400





Wages 3254





Indirect Business Taxes 500





Interest 530





Domestic Investment 741





Government Expenditure 1098





Rental Income 17





Corporate Profits 744





Exports 673





Imports 704





Net Factor Payment 20






1
Expert's answer
2022-02-01T11:14:42-0500

Y = C + I + G + NX

Y = 3,657 + 741 + 1,098 + (673 – 704)

Y = 3,657 + 741 + 1,098 – 31

Y = 5,465